# Position Coupling: Leveraging Task Structure for Improved Length Generalization of Transformers

### v1

@@ A.2 Two-dimensional Task: Sharing the Position Embedding Modules or Not
Figure 18: Exact-match accuracy (median over 4 runs) for minesweeper generator task, trained on sequences of length (5–9)$\times$(5–9) with position coupling. For further experiment details, refer to the Table 8.
Appendix B Experiment Details and Hyperparameters
Position coupling can be easily implemented on top of usual libraries of training transformer models like HuggingFace (Wolf et al., 2019) and Flaxformer555github.com/google/flaxformer since these libraries support an arbitrary array of position IDs (in the case of using APE). All we need is to build up a short routine implementing the assigning rule of position IDs when establishing the dataset and data loaders. To compare with NoPE, we use the code base provided by Kazemnejad et al. (2023) for most of the experiments.666github.com/McGill-NLP/length-generalization It contains a custom implementation of decoder-only T5 (Raffel et al., 2020) established on top of PyTorch (Paszke et al., 2019) and Huggingface, including several PE methods. We additionally implement a custom RMSNorm module (Zhang and Sennrich, 2019) and various positioning schemes of normalization layers (e.g., PreNorm (Xiong et al., 2020), PostNorm (Vaswani et al., 2017), and their combination), to follow the implementation details of Zhou et al. (2024b).
Table 1: Hyperparameter summary for decimal integer addition task in the main text: comparison between trained lengths (Figure 4).
@@ Appendix B Experiment Details and Hyperparameters
Table 2: Hyperparameter summary for decimal integer addition task in the main text: comparison between the number of layers (Figure 5).
@@
Embedding Dimension 512
Dimension per Head 64
@@
Trained Lengths of Operands 1–40
@@
Table 7: Hyperparameter summary for copy/reverse task (Figures 12 and 13).
@@
Table 8: Hyperparameter summary for minesweeper generator task (Figures 15 and 18).
@@
Appendix C Decoder-only Transformer Architecture
Here we detail the architecture of an depth-$L$, $H$-head decoder-only Transformer (Vaswani et al., 2017). For a simple presentation, we ignore the normalization layers, as in Yun et al. (2020a).
@@ Appendix C Decoder-only Transformer Architecture
Appendix D Formal Construction of Addition Transformer with Position Coupling
Here we show how to implement the addition by employing a depth-1 two-head decoder-only Transformer equipped with position coupling. We restate the theorem for the sake of readability.
@@ Appendix D Formal Construction of Addition Transformer with Position Coupling
  • We start with the notation (Section D.1).
  • We review and formalize the format of the input sequence (zero-padding, reversed format, and wrapping with BOS/EOS) (Section D.2).
  • We define the encoding function ${\tt Enc}$ with a table of a concrete example (Section D.3), where ${\tt Enc}$ maps an input sequence of length $N$ to a $d\times N$ encoding matrix ${\bm{X}}^{(0)}$.
  • We devote a lot of pages to the construction of the parameters of a causal attention layer ${\tt Att}_{1}$ to generate desired attention patterns (Section D.4). The attention layer has two heads playing distinct roles: (1) preparing for a sum without considering carries; and (2) preparing for the carry prediction & EOS detection.
  • We provide a construction of a token-wise feed-forward neural network ${\tt FF}_{1}$ which is a two-layer ReLU network (Section D.5). It consists of two subnetworks playing different roles: (1) producing one-hot vectors, each of which indicates a digit of the sum (response); and (2) binary values indicating whether the position is the end of the sequence.
  • We conclude the proof by defining the decoding function ${\tt Dec}$ which performs the linear readout and the arg-max operation to generate the output sequence (Section D.6).
@@
D.1 Notation
For the architecture of the decoder-only Transformer, we follow the notation introduced in Appendix C.
@@ D.2 Input Sequence
We first zero-pad the shorter one between $a$ and $b$ to match the length of the part $A$ and part $B$ as $\ell$. Sometimes, the sum $c$ might be longer than $a$ or $b$ due to a carry. To make the length of the part $C$ consistent, we also put a zero-pad in front of $c$ to set its length as $\ell+1$. Also, to ease calculating the addition with next-token prediction, we reverse the sum $c$ to make the part $C$. For example, if we have a sum $3812+98=3910$, we use \$$\overline{\$3812+{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}00}98={\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}0193}{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}0}}$ as an input sequence; if a sum $98+9907=10005$ is given, we use \$$\overline{\${\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}00}98+9907={\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}50001}}$ as an input sequence. The red digits are zero-paddings, and the blue digits are the reversed sum.
@@ D.3.1 Token Embedding
Dimension 1 (num).   For a number token ($0,\ldots,9$), we put itself in the dimension num. For the other tokens (+,=,\$currency-dollar+,=,\\$+ , = , \$), we put 00.
Dimension 2 (is_bos).   For a special token ‘\$currency-dollar\\$\$’, we put $1$ in the dimension is_bos. We put 00 otherwise.
Dimension 3 (full_ones).   We put $1$ everywhere in this dimension.
D.3.2 Coupled Position IDs and Position Embedding
@@
D.4 Transformer Block — Causal Attention Layer
@@
D.4.1 Attention Head 1: Digit-wise Addition without Carries
The goal of the first head is to perform a digit-wise addition without carries and to fill in the blanks of the encoding matrix at dimension pre_sum. Later, using this dimension, combined with the dimension pre_carry, we will be able to perform the next-token prediction for addition. For now, we do not care about the carries, which will be dealt with in a later section. Formally, we aim to perform $\sigma_{i}+\sigma_{i+\ell+1}$ for each $i\in\{2,\cdots,\ell+1\}$ and put its result at the $(3\ell+4-i)$-th position (column) of the dimension pre_sum (row). To this end, we utilize our position embedding.
@@ D.4.1 Attention Head 1: Digit-wise Addition without Carries
D.4.2 Attention Head 2: Carry & EOS Detection
@@ D.4.3 Residual Connection
$(P+18)$$(2P+17)$: pos_2 ${\bm{0}}_{P}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{3}$
D.5 Transformer Block — Token-wise Feed-forward Layer
@@
Recall from the discussion in Section D.4.2 that we can judge whether a carry 1 is generated at a certain position by exploiting only the digits (of the operands and the sum) in the same significance. Bringing the notation, let $a$ and $b$ be digits of the operands in the same significance, and $c$ be a digit of the sum in the same significance as $a$ and $b$. Then the rule of recognizing that the addition of $a$ and $b$ generates a carry is that
@@
D.5.1 Subnetwork 1: Construction for sum (dimension 7–16).
@@
Then we construct the first subnetwork of our feed-forward network by
$\left[{\tt FF}_{1}\left({\bm{Y}}^{(1)}\right)\right]_{(\textsc{sum})j}=\begin{bmatrix}f_{0}\left(g\left({\bm{Y}}^{(1)}_{\bullet j}\right)\right)&\cdots&f_{9}\left(g\left({\bm{Y}}^{(1)}_{\bullet j}\right)\right)\end{bmatrix}^{\top}.\quad(j\in[N])$ (60)
@@
D.5.2 Subnetwork 2: Construction for is_eos (dimension 17).
@@
Then, we construct the second subnetwork of our feed-forward network by
$\left[{\tt FF}_{1}\left({\bm{Y}}^{(1)}\right)\right]_{(\textsc{is\_eos})j}=h\left({\bm{Y}}^{(1)}_{(\textsc{pre\_eos})j}\right).\quad(j\in[N])$ (68)
@@
Remark:
  • In total, we consume $8\times 10+2=82$ ReLU neurons in our feed-forward network ${\tt FF}_{1}$. However, it is possible to construct the addition Transformer with a smaller number of neurons, with a slight modification in the linear readout of the decoding function (Section D.6).
@@ D.5.3 Residual Connection
$(P+18)$$(2P+17)$: pos_2 ${\bm{0}}_{P}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{3}$
D.6 Decoding Function
As mentioned in Appendix C, the decoding function performs a linear readout (with a weight matrix ${\bm{W}}_{\rm out}\in\mathbb{R}^{\left|{\mathcal{V}}\right|\times d}$) and a (token-wise) arg-max operation. That is,
@@ D.6 Decoding Function
Table 26: Example output of linear readout (${\bm{W}}_{\rm out}{\bm{X}}^{(1)}$), continuing from Tables 24 and 25. The yellow cells represent the maximum value of each column, from the ‘=’ token’s column to the rightmost column (which are used for next-token prediction).
@@
Appendix E Impossibility of Addition with No Positional Encoding
@@
We assume the 1-layer ($L=1$) $H$-head Transformer achitecture specified in Appendix C. Although it omits normalization layers, we remark that Proposition 5.2 remains valid even for the architecture with a standard layer normalization (Ba et al., 2016) or its variants (e.g., Zhang and Sennrich, 2019).
@@
We keep following the notation about matrices introduced in Section D.1. Throughout the proof, we denote the value/vector/matrix related to ${\mathcal{I}}^{\prime}$ by appending ‘’ to it.
@@
The claim suffices to prove the proposition because of the following: first, the claim implies that the last ($N$-th) column of the attention layer outputs are the same, i.e., $\left[{\tt Att}({\bm{X}})\right]_{\bullet N}=\left[{\tt Att}({\bm{X}}^{\prime})\right]_{\bullet N}$. Note that the operations after the attention layer—residual connections, ${\tt FF}$, and ${\tt Dec}$—all operate in a token-wise (column-by-column) manner: the $j$-th column of the output of a token-wise operation is a function of $j$-th column of the input for the operation. Therefore, the last column of the attention layer output totally determines the next-token prediction at $N$-th input token. As a result, the predicted next-tokens are the same for ${\mathcal{I}}$ and ${\mathcal{I}}^{\prime}$.
The rest of the proof is devoted to prove the aforementioned claim. Fix any $h\in[H]$. Let
@@ 48 Total m-digit Additions Count: 8100000000
Appendix F (Formal) Construction of $N\times 2$ Multiplication Transformer with Position Coupling
@@ F.2 Input Sequence
Unlike the addition case, we do not zero-pad both $a$ and $b$. Instead, we only zero-pad the response, as the length of $c$ may either equal the sum of the lengths of $a$ and $b$, or be less than the sum of their lengths by $1$. Hence, we zero-pad in front of $c$ for the latter case to fix the length of $c$ by $\ell$. We also reverse the response $c$ to make the part $C$. For instance, if we have $312\times 24=7488$, the input sequence transforms to \$$\overline{\$312\times 24={\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}8847}{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}0}}$. If we have $589\times 62=36518$, then the input sequence would be \$$\overline{\$589\times 62={\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}81563}}$. The red digit is a zero-padding, and the blue digits are the reversed product.

### v2

@@ B.8 Position Coupling for Minesweeper Generator Tasks
Figure 21: Exact-match accuracy (median over 4 runs) for minesweeper generator task, trained on sequences of length (5–9)$\times$(5–9) with position coupling. For further experiment details, see Table 8.
Appendix C Experiment Details and Hyperparameters
Position coupling can be easily implemented on top of usual libraries of training transformer models like HuggingFace (Wolf et al., 2019) and Flaxformer444github.com/google/flaxformer since these libraries support an arbitrary array of position IDs (in the case of using APE). All we need is to build up a short routine implementing the assigning rule of position IDs when establishing the dataset and data loaders. To compare with NoPE, we use the code base provided by Kazemnejad et al. (2023) for most of the experiments.555github.com/McGill-NLP/length-generalization It contains a custom implementation of decoder-only T5 (Raffel et al., 2020) established on top of PyTorch (Paszke et al., 2019) and Huggingface, including several PE methods. We additionally implement a custom RMSNorm module (Zhang and Sennrich, 2019) and various positioning schemes of normalization layers (e.g., PreNorm (Xiong et al., 2020), PostNorm (Vaswani et al., 2017), and their combination), to follow the implementation details of Zhou et al. (2024b).
Table 1: Hyperparameter summary for decimal integer addition task: comparison between trained lengths (Figures 3 and 14).
@@ Appendix C Experiment Details and Hyperparameters
Table 2: Hyperparameter summary for decimal integer addition task: comparison between the number of layers (Figures 4 and 15).
@@
Embedding Dimension 1024
Dimension per Head 128
@@
Trained Lengths of Operands 1–30
@@
Table 7: Hyperparameter summary for copy/reverse task (Figure 20).
@@
Table 8: Hyperparameter summary for minesweeper generator task (Figures 10 and 21).
@@
Appendix D Decoder-only Transformer Architecture
Here we detail the architecture of a depth-$L$, $H$-head decoder-only Transformer (Vaswani et al., 2017). For a simple presentation, we ignore the normalization layers, as in Yun et al. (2020a).
@@ Appendix D Decoder-only Transformer Architecture
Appendix E Formal Construction of Addition Transformer with Position Coupling
Here we show how to implement the addition by employing a single-layer two-head decoder-only Transformer equipped with position coupling. We restate the theorem for the sake of readability.
@@ Appendix E Formal Construction of Addition Transformer with Position Coupling
  • We start with the notation (Section E.1).
  • We review and formalize the format of the input sequence (zero-padding, reversed format, and wrapping with BOS/EOS) (Section E.2).
  • We define the encoding function ${\tt Enc}$ with a table of a concrete example (Section E.3), where ${\tt Enc}$ maps an input sequence of length $N$ to a $d\times N$ encoding matrix ${\bm{X}}^{(0)}$.
  • We devote a lot of pages to the detailed construction of the parameters of a causal attention layer ${\tt Att}_{1}$ to generate desired attention patterns (Section E.4). The attention layer has two attention heads playing distinct roles: (1) preparing for a sum without considering carries; and (2) preparing for the carry prediction & EOS detection.
  • We provide a construction of a token-wise feed-forward neural network ${\tt FF}_{1}$ which is a two-layer ReLU network (Section E.5). It consists of two subnetworks playing different roles: (1) producing one-hot vectors, each of which indicates a digit of the sum (response); and (2) binary values indicating whether the position is the end of the sequence.
  • We conclude the proof by defining the decoding function ${\tt Dec}$ which performs the linear readout and the arg-max operation to generate the output sequence (Section E.6).
@@
E.1 Notation
For the architecture of the decoder-only Transformer, we follow the notation introduced in Appendix D.
@@ E.2 Input Sequence
We first zero-pad the shorter one between $a$ and $b$ to match the length of the part $A$ and part $B$ as $\ell$. Sometimes, the sum $c$ might be longer than $a$ or $b$ due to a carry. To make the length of the part $C$ consistent, we also put a zero-pad in front of $c$ to set its length as $\ell+1$. Also, to ease calculating the addition with next-token prediction, we reverse the sum $c$ to make the part $C$. For example, if we have a sum $3812+98=3910$, we use \$$\overline{\$3812+{\color[rgb]{.804,.15,.15}00}98={\color[rgb]{.094,.455,.804}0193}{\color[rgb]{.804,.15,.15}0}}$ as an input sequence; if a sum $98+9907=10005$ is given, we use \$$\overline{\${\color[rgb]{.804,.15,.15}00}98+9907={\color[rgb]{.094,.455,.804}50001}}$ as an input sequence. The red digits are zero-paddings, and the blue digits are the reversed sum.
@@ E.3.1 Token Embedding
Dimension 1 (num).
For a number token ($0,\ldots,9$), we put itself in the dimension num. For the other tokens (+,=,\$currency-dollar+,=,\\$+ , = , \$), we put 00.
Dimension 2 (is_bos).
For a special token ‘\$currency-dollar\\$\$’, we put $1$ in the dimension is_bos. Otherwise, we put 00.
Dimension 3 (full_ones).
We put $1$ everywhere in this dimension.
E.3.2 Coupled Position IDs and Position Embedding
@@
E.4 Transformer Block — Causal Attention Layer
@@
E.4.1 Attention Head 1: Digit-wise Addition without Carries
The goal of the first head is to perform a digit-wise addition and to fill in the blanks of the encoding matrix at dimension pre_sum. Later, using this dimension, combined with the dimension pre_carry, we will be able to perform the next-token prediction for addition. For now, we do not care about the carries, which will be dealt with in a later section. Formally, we aim to perform $\sigma_{i}+\sigma_{i+\ell+1}$ for each $i\in\{2,\cdots,\ell+1\}$ and put its result at the $(3\ell+4-i)$-th position (column) of the dimension pre_sum (row). To this end, we utilize our position embedding.
@@ E.4.1 Attention Head 1: Digit-wise Addition without Carries
E.4.2 Attention Head 2: Carry & EOS Detection
@@ E.4.3 Residual Connection
$(P+18)$-$(2P+17)$: pos_2 ${\bm{0}}_{P}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{3}$
E.5 Transformer Block — Token-wise Feed-forward Layer
@@
Recall from the discussion in Section E.4.2 that we can judge whether a carry 1 is generated at a certain position by exploiting only the digits (of the operands and the sum) in the same significance. Bringing the notation, let $a$ and $b$ be digits of the operands in the same significance, and $c$ be a digit of the sum in the same significance as $a$ and $b$. Then the rule of recognizing that the addition of $a$ and $b$ generates a carry is that
@@
E.5.1 Subnetwork 1: Construction for sum (dimension 7–16).
@@
Then we construct a subnetwork of our feed-forward network for a token $\sigma_{j}$ by
$\left[{\tt FF}_{1}\left({\bm{Y}}^{(1)}\right)\right]_{(\textsc{sum})j}=\begin{bmatrix}f_{0}\left(g\left({\bm{Y}}^{(1)}_{\bullet j}\right)\right)&\cdots&f_{9}\left(g\left({\bm{Y}}^{(1)}_{\bullet j}\right)\right)\end{bmatrix}^{\top}.$ (60)
@@
E.5.2 Subnetwork 2: Construction for is_eos (dimension 17).
@@
Then, we can construct a subnetwork of our feed-forward network for a token $\sigma_{j}$ by
$\left[{\tt FF}_{1}\left({\bm{Y}}^{(1)}\right)\right]_{(\textsc{is\_eos})j}=h\left({\bm{Y}}^{(1)}_{(\textsc{pre\_eos})j}\right).$ (68)
@@
Remarks:
  • In total, we consume $8\times 10+2=82$ ReLU neurons in our feed-forward network ${\tt FF}_{1}$. However, it is possible to construct the addition Transformer with a smaller number of neurons, with a slight modification in the linear readout of the decoding function (Section E.6).
@@ E.5.3 Residual Connection
$(P+18)$-$(2P+17)$: pos_2 ${\bm{0}}_{P}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{7}$ ${\bm{v}}^{P}_{6}$ ${\bm{v}}^{P}_{5}$ ${\bm{v}}^{P}_{4}$ ${\bm{v}}^{P}_{3}$
E.6 Decoding Function
As mentioned in Appendix D, the decoding function performs a linear readout (with a weight matrix ${\bm{W}}_{\rm out}\in\mathbb{R}^{\left|{\mathcal{V}}\right|\times d}$) and a (token-wise) arg-max operation. That is,
@@ E.6 Decoding Function
Table 26: Example output of linear readout (${\bm{W}}_{\rm out}{\bm{X}}^{(1)}$), continuing from Tables 24 and 25. The yellow cells represent the maximum value of each column, from the ‘=’ token’s column to the rightmost column (used for next-token prediction).
@@
Appendix F Impossibility of Addition with No Positional Encoding
@@
We assume the 1-layer ($L=1$) $H$-head Transformer achitecture specified in Appendix D. Although it omits normalization layers, we remark that Proposition 5.2 remains valid even for the architecture with a standard layer normalization (Ba et al., 2016) or its variants (e.g., Zhang and Sennrich, 2019).
@@
We keep following the notation about matrices introduced in Section E.1. Throughout the proof, we denote the value/vector/matrix related to ${\mathcal{I}}^{\prime}$ by appending ‘’ to it.
@@
The claim suffices to prove the proposition because of the following: first, the claim implies that the last ($N$-th) columns of the attention layer outputs are the same, i.e., $\left[{\tt Att}({\bm{X}})\right]_{\bullet N}=\left[{\tt Att}({\bm{X}}^{\prime})\right]_{\bullet N}$. Note that the operations after the attention layer—residual connections, ${\tt FF}$, and ${\tt Dec}$—all operate in a token-wise (column-by-column) manner: the $j$-th column of the output of a token-wise operation is a function of $j$-th column of the input for the operation. Therefore, the last column of the attention layer output totally determines the next-token prediction at $N$-th input token. As a result, the predicted next-tokens are the same for ${\mathcal{I}}$ and ${\mathcal{I}}^{\prime}$.
The rest of the proof is devoted to proving the aforementioned claim. Fix any $h\in[H]$. Let
@@ 48 Total m-digit Additions Count: 8100000000
Appendix G (Formal) Construction of $N\times 2$ Multiplication Transformer with Position Coupling
@@ G.2 Input Sequence
Unlike the addition case, we do not zero-pad both $a$ and $b$. Instead, we only zero-pad the response, as the length of $c$ may either equal the sum of the lengths of $a$ and $b$, or be less than the sum of their lengths by $1$. Hence, we zero-pad in front of $c$ for the latter case to fix the length of $c$ by $\ell$. We also reverse the response $c$ to make the part $C$. For instance, if we have $312\times 24=7488$, the input sequence transforms to \$$\overline{\$312\times 24={\color[rgb]{.094,.455,.804}8847}{\color[rgb]{.804,.15,.15}0}}$. If we have $589\times 62=36518$, then the input sequence would be \$$\overline{\$589\times 62={\color[rgb]{.094,.455,.804}81563}}$. The red digit is a zero-padding, and the blue digits are the reversed product.